OpenStudy (anonymous):

(x+2)/(2x)=(1)/(x) simplify rational equations by cross multiplying and factoring

OpenStudy (anonymous):

@ganeshie8 @zepdrix

OpenStudy (anonymous):

ok cross multiply \[x\times (x+2)=1\times 2x\]

OpenStudy (anonymous):

Now what do we do? Im not so great at this and dont really understand it :/ @satellite73

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@Imtiaz7 Do you have an idea about this?

OpenStudy (imtiaz7):

yes i have but first Lissie u should multiply (x+2) by x and then 2x multiply by 1 so we get x^2+2x=2x x^2=0 bc 2x are cancelled with each other so finally we get x=0

OpenStudy (anonymous):

Okay, Thank you and so then to solve we just put in 0 for x right?

OpenStudy (imtiaz7):